OpenStudy (anonymous):

An unknown gas effuses at .850 times the effusion rate of NO2. Estimate the molar mass of the unknown gas.

OpenStudy (anonymous):

\[\frac{ r1 }{ r2 } = \sqrt{\frac{ M2 }{ M1 }}\]

OpenStudy (anonymous):

r1 = 0.85r2 M2 = 14 + 32 = 46 Find M1...thats all